What is the limiting reactant if you have 100 grams of NaOH and 100 grams of Al?
6NaOH + 2Al ---> 2Na3AlO3 + 3H2
1
Expert's answer
2017-03-02T05:42:23-0500
In a chemical reaction, the limiting reagent, also known as the "limiting reactant", is the substance which is totally consumed when the chemical reaction is complete. The amount of product formed is limited by this reagent since the reaction cannot proceed further without it. The other reagents may be present in excess of the quantities required to react with the limiting reagent.

For answering this question you need to find amount of every reactant.

n=m/Mw, where Mw is molecular weight of compound
n of NaOH is 100/40 = 2.5 mol
n of Al is 100/27=3.7 mol
The ratio of amounts of NaOH : Al, must be 6:3 or 2: 1.
2.5 mol of NaOH can react with 2.5/2= 1.25 mol of Al, but you have 3.7,so NaOH is the limiting reagent.